# WaterMax: breaking the LLM watermark detectability-robustness-quality trade-off

### v1

@@ Abstract
LLM, Watermarking
@@ 1 Introduction
The availability of powerful large-language models (LLMs) allows users to produce texts that look like human writings. The risk for misuse of these models is critical, ranging from the impersonation of individuals to the large-scale generation of fake news. Identifying the provenance of a given piece of text is paramount to limit the impact of such ‘weaponization’ of LLMs. New initiatives or regulations impose technical means for AI traceability (NIST, 2023; EUParliament, 2023; CAC, 2022). There exists two main classes of methods to perform such identification: passive methods,   i.e., forensics, and active methods,   i.e., watermarking. Forensics methods for generated texts generally leverage a priori knowledge about the statistics of texts generated by a given class of LLMs (Wu et al., 2023; Mitchell et al., 2023). Despite their versatility, these methods offer lower performance than watermarking and do not provide strong guarantees on false alarm rates. The reported probabilities are only validated empirically on some datasets, and because of this, they are never lower than $10^{-3}$ (Hans et al., 2024).
Figure 1: The relative quality of the watermarked text as a function of the watermark size (i.e. the number of tokens necessary to detect the watermark - see definitions in Sect. 7). Transparency indicates a less robust scheme. Our watermark attains a small watermark size without loosing quality. [^1] The figure is adapted from Fig. 5 of (Piet et al., 2023) with values graciously shared by the authors – except for our scheme..
@@
In contrast, active methods are only limited by the fact the LLM owner must integrate the watermarking within the generation processes. This is done by embedding an imperceptible signal in the generated text, which can be retrieved by a detector sharing the secret key of the model owner. Current watermarking methods for generative texts (Kirchenbauer et al., 2023b; Aaronson & Kirchner, 2023; Kuditipudi et al., 2023; Christ et al., 2023) have shown to lead to extremely high detection rates with an acceptable impact on text quality. Despite that, they are all prone to the same limitations described in Section 2. This paper presents WaterMax, a new watermarking technique for LLM-generated text. It obeys the regular constraints found in the literature: it can be integrated into any standard LLM without fine-tuning the weights, the detection does not need the original LLM, and it can spot the watermark on a slice of generated text with some guarantee on the false alarm rate. Our contributions are the following:
  • Our scheme improves a lot the quality of the generated text as well as the detectability of the watermark. Fig. 1 clearly shows that the other methods need to boost the watermark strength to become detectable on short texts, inevitably degrading the quality, whereas our scheme sustains a pristine quality of the generated text.
@@
  • Our design is new. The high text quality comes from the fact that it does not rely on the usual mechanisms of the literature, especially it keeps the next token distribution, the sampling temperature and technique intact. Moreover, it better utilizes the text entropy by working over sub-sequences of tokens, so-called chunks hereafter, rather than token by token.
@@
  • Our study provides a theoretical model of the watermark performance characterizing the false positive and the true positive rates even under attack.
@@ 2 Related Work
Watermarking texts generated by LLMs is mainly performed by changing either the sampling distribution of tokens (Kirchenbauer et al., 2023a) or directly the method of selecting the next token (Aaronson & Kirchner, 2023; Kuditipudi et al., 2023; Christ et al., 2023). Critically, the false-positive rate of these methods can be reliably controlled (Fernandez et al., 2023). Furthermore, the computational cost of most of these methods is negligible relative to the text generation itself.
On the other hand, they all suffer from similar weaknesses:
Token entropy limit
Recent theoretical works (Christ et al., 2023; Huang et al., 2023; Kirchenbauer et al., 2023a; Aaronson & Kirchner, 2023) show that the watermark detectability is highly dependent on the entropy of the next token distribution at each step. In practice, this makes the text length necessary for a reliable watermark detectability dependent on the type of text. One solution is to adapt the temperature in the LLM to increase the entropy artificially (Idrissi et al., 2023; Liu & Bu, 2024).
Quality of the generated text
Kuditipudi et al. (2023) claim that a watermark is distortion-free provided it does not modify the probability distribution of the next token on average over the keys. This is guaranteed, for instance, in schemes like (Aaronson & Kirchner, 2023; Kuditipudi et al., 2023; Christ et al., 2023). However, as shown in (Piet et al., 2023), this does not lead to a better watermarking scheme in practice, as individual texts still incur a non-negligible distortion. Other techniques like (Kirchenbauer et al., 2023a) modify the probability distribution of the next token, and it is not straightforward to predict the impact on the quality. In practice, there is a trade-off between text quality and watermark strength, with more detectable watermarks leading to significant text degradation (Fernandez et al., 2023; Piet et al., 2023).
Text size limit
All these schemes are based on a score function that accumulates watermarking evidences over individual tokens. Consequently, their power is an increasing function of the text size, with small texts being more difficult to watermark without losing quality.
Our new watermarking method addresses all these problems. In particular, our scheme can theoretically reach an arbitrary watermarking power without any loss in text quality, even for short texts given enough entropy. This is done by directly minimizing the $p$-value of a full text for a watermark detector instead of performing an ad-hoc manipulation of individual tokens. This method has a computational cost, which we show how to limit throughout the paper.

### v2

@@ Abstract
Figure 1: Detectability as a function of text quality for different LLM architectures. WaterMax always reaches a detectability close to $1$ despite a negligible loss of quality. Probability of false-alarm fixed at $10^{-6}$, nucleus sampling ($top_{p}=0.95$) at temperature $1.0$. Text quality is measured as the relative perplexity of the watermarked text over the non-watermarked text.
@@ 1 Introduction
The availability of powerful large-language models (LLMs) allows users to produce texts that look like human writings. The risk for misuse of these models is critical, ranging from the impersonation of individuals to the large-scale generation of fake news. Identifying the provenance of a given piece of text is paramount to limit the impact of such ‘weaponization’ of LLMs. New initiatives or regulations impose technical means for AI traceability [27, 9, 5].
Forensics passive methods generally leverage a priori knowledge about the statistics of texts generated by a given class of LLMs [33, 26]. Despite their versatility, these methods offer low performance. The reported probabilities of errors are only validated empirically on some datasets, and because of this, they are never lower than $10^{-3}$ [14].
In contrast, active methods like watermarking are only limited by the fact the LLM owner must integrate the watermarking within the generation processes. This is done by embedding an imperceptible signal in the generated text, which can be retrieved by a detector sharing the secret key of the model owner. Current watermarking methods for generative texts [21, 1, 22, 8] provide low and guaranteed false positives rates. Yet, the trade-off between the detectability and the text quality crucially depends on the entropy of the text to be generated, which in turn depends on the prompt and the LLM, as illustrated in Fig. 1. This implies that the distortion-free property [1, 22, 8] ensures that watermarking does not degrade text quality but inherently limits the detectability.
@@
This paper presents WaterMax, a watermarking technique that trades off robustness not for quality, but for complexity. It obeys the regular constraints found in the literature: it can be integrated into any standard LLM without fine-tuning the weights, the detection does not need the original LLM, and it can spot the watermark on a slice of generated text with some guarantee on the false positive rate. Our contributions are the following:
@@
  • WaterMax is based on a new design not relying on the usual mechanisms of the literature; especially, it keeps the next token distribution and sampling (temperature and method) intact. Moreover, it better utilizes the text entropy by working over sub-sequences of tokens, so-called chunks hereafter, rather than token by token.
  • This new design makes WaterMax almost distortion-free, as shown experimentally (Sect. 7.3) and justified theoretically (App. H), and yet enjoys higher robustness than the state-of-the-art, consistently over several LLMs. Figure 1 shows that the other methods need to boost the watermark strength to be as detectable as WaterMax, inevitably degrading the quality.
@@
  • This new design facilitates building up a theoretical model of the watermark robustness characterizing the true positive rates under attack (see Prop. 5.2).
@@ 2 Related Work
Watermarking texts generated by LLMs is mainly performed by changing either the distribution [20] or the sampling [1, 22, 8] of the next token. Critically, the false-positive rate of these methods can be reliably controlled [11]. Furthermore, the computational cost of most of these methods is negligible relative to the text generation itself. On the other hand, they all suffer from similar weaknesses:
Text entropy limit
Some theoretical works [8, 17, 20, 1] show that the watermark detectability is highly dependent on the entropy of the text to be generated. In practice, this makes the text length necessary for reliable watermark detectability dependent on the type of the text, and also the LLM. One may increase the watermark strength or the temperature in the LLM to increase the entropy artificially [19, 25]. Both solutions degrade text quality compared to the original LLM.
Distortion-freeness and quality
Kuditipudi et al. [22] define that a watermark is distortion-free if it does not modify the probability distribution of the next token on average over the keys. This is guaranteed in schemes like [1, 22, 8]. Yet, these schemes rely heavily on the entropy of the token distribution, leading to possibly low detectability without any means to increase it without losing distortion-freeness. Appendix J illustrates the dependence of Aaronson’s scheme on the original LLM. Moreover, a scheme that is not distortion-free does not necessarily lead to texts with lower empirical quality, but foreseeing the impact of the watermark strength on the quality is an issue.